OpenStudy (mew55):

one last question integral 0 to infinity dx/sqrt[x]

OpenStudy (mew55):

\[\int\limits_{0}^{\infty} \frac{ dx }{ \sqrt{x} }\]

OpenStudy (cwrw238):

= INT x^(-0.5) = x^(0.5) / 0.5 = 2 x^(0.5)

OpenStudy (mew55):

u use the \[\frac{ x^(n+1) }{ n+1 } \] right?

OpenStudy (cwrw238):

yes - thats what i did you can also write it as 2 sqrt x

OpenStudy (mew55):

then take the limits of 0 to infinity right?

OpenStudy (cwrw238):

between the limits 0 and infinity its value is infinity - 0 = infinity

OpenStudy (mew55):

which is diverges right?

OpenStudy (cwrw238):

yep

OpenStudy (mew55):

kk :3. thank u. gosh i hate when those problems r that simple yet i made it complicated :S

OpenStudy (cwrw238):

lol - keep practicing

OpenStudy (mew55):

im hanging in that cal II class. got to get my gpa up back to .3 or 1.3 points